### v1

# Analysis of Emotion in Rumour Threads on Social Media

Abstract
Rumours in online social media pose significant risks to modern society, motivating the need for better understanding of how they develop. We focus specifically on the interface between emotion and rumours in threaded discourses, building on the surprisingly sparse literature on the topic which has largely focused on emotions within the original rumour posts themselves, and largely overlooked the comparative differences between rumours and non-rumours. In this work, we provide a comprehensive analytical emotion framework, contrasting rumour and non-rumour cases using existing NLP datasets to further understand the emotion dynamics within rumours. Our framework reveals several findings: rumours exhibit more negative sentiment and emotions, including anger, fear and pessimism, while non-rumours evoke more positive emotions; emotions are contagious in online interactions, with rumours facilitate negative emotions and non-rumours foster positive emotions; and based on causal analysis, surprise acts as a bridge between rumours and other emotions, pessimism is driven by sadness and fear, optimism by joy and love.
Introduction
In part due to the ubiquity of edge devices like mobile phones, the major of the world’s population now has access to the internet. [^1] https://datareportal.com/reports/digital-2024-october-global-statshot This increasing ease of access and interaction through online social media has brought both opportunities and challenges. One significant challenge is the rapid spread of rumours. Rumours on online social media have become a major threat to society (Tian, Zhang, and Lau 2022; Zubiaga et al. 2015; Kochkina, Liakata, and Zubiaga 2018; Ma, Gao, and Wong 2017). The circulation of unsubstantiated rumours has impacted a large group of people, with consequences ranging from seeding skepticism and discrediting science, to endangering public health and safety. For example, during COVID-19, an Arizona man died, and his wife was hospitalized after ingesting a form of chloroquine in an attempt to prevent the disease. Additionally, 77 cell phone towers were set on fire due to conspiracy theories linking 5G networks to the spread of COVID-19 (Cui and Lee 2020). Recent advancements in Large Language Models (LLM) and generative AI have exacerbated this phenomenon, creating an urgent need to understand and better deal with rumours on social media (Chen and Shu 2024).
Previous research has highlighted several factors driving the spread of rumours on social media (Zollo et al. 2015). These factors often relate to the characteristics of publishers; for instance, users with more followers can reach wider audiences, and the number of reshares and likes reflects users’ beliefs and attitudes toward a post (Zaman, Fox, and Bradlow 2013; Vosoughi, Roy, and Aral 2018). Other studies have focused on the online diffusion of specific topics, such as elections or disasters (Starbird 2017; Domenico et al. 2013), and other harmful online social contents (Aleksandric et al. 2024).
Emotions have a strong influence on human behavior in both offline and online settings (Zollo et al. 2015; Herrando and Constantinides 2021; Ekman 1992). They shape the type of information users seek, how they process and remember it, and the judgments and decisions they make. Misinformation is often associated with high-arousal emotions such as anger, sadness, anxiety, surprise, and fear (Liu et al. 2024b). Rumours conveying these emotions are more likely to generate higher numbers of shares and exhibit long-lived, viral patterns (Pröllochs, Bär, and Feuerriegel 2021b). However, existing research on emotions in misinformation analysis is fragmented, focusing primarily on the emotions within the original rumour posts themselves and often overlooking the comparative differences between rumours and non-rumours (Ferrara and Yang 2015; Zollo et al. 2015). In this work, we address this gap by introducing a systematic analysis of various aspects of emotion, contrasting the emotional patterns in rumour and non-rumour content on social media. Using popular NLP rumour detection datasets, we provide new insights into the emotional dynamics of rumour and non-rumour data.
Data
We use three rumour datasets in this work, namely: PHEME (Zubiaga et al. 2015; Kochkina, Liakata, and Zubiaga 2018), Twitter15, and Twitter16 (Ma, Gao, and Wong 2017):
PHEME
Zubiaga et al. (2015) contains 6,425 tweet posts of rumours and non-rumours related to 9 events. To avoid using specific a priori keywords to search for tweet posts, PHEME used the Twitter (now X) steaming API to identify newsworthy events from breaking news and then selected from candidate rumours that met rumour criteria, finally they collected associated conversations and annotate them. They engaged journalists to annotate the threads. The data were collected between 2014 and 2015. The 9 events are split into two groups, the first being breaking news that contains rumours, including Ferguson unrest, Ottawa shooting, Sydney siege, Charlie Hebdo shooting, and Germanwings plane crash. The rest are specific rumours, namely Prince to play in Toronto, Gurlitt collection, Putin missing, and Michael Essien contracting Ebola.
Twitter 15
Liu et al. (2015) was constructed by collecting rumour and non-rumour posts from the tracking websites snopes.com and emergent.info. They then used the Twitter API to gather corresponding posts, resulting in 94 true and 446 false posts. This dataset further includes 1,490 root posts and their follow posts, comprising 1,116 rumours and 374 non-rumours.
Twitter 16
Similarly to Twitter 15, Ma et al. (2016) collected rumours and non-rumours from snopes.com, resulting in 778 reported events, 64% of which are rumours. For each event, keywords were extracted from the final part of the Snopes URL and refined manually—adding, deleting, or replacing words iteratively—until the composed queries yielded precise Twitter search results. The final dataset includes 1,490 root tweet posts and their follow posts, comprising 613 rumours and 205 non-rumours.
Task Prompt
V-oc Categorize the text into an ordinal class that best characterizes the writer’s mental state, considering various degrees of positive and negative sentiment intensity. 3: very positive mental state can be inferred. 2: moderately positive mental state can be inferred. 1: slightly positive mental state can be inferred. 0: neutral or mixed mental state can be inferred. -1: slightly negative mental state can be inferred. -2: moderately negative mental state can be inferred. -3: very negative mental state can be inferred.
E-c Categorize the text’s emotional tone as either ‘neutral or no emotion’ or identify the presence of one or more of the given emotions (anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust).
E-i Assign a numerical value between 0 (least E) and 1 (most E) to represent the intensity of emotion E expressed in the text.
Table 1: Prompts used for EmoLLM to detect emotion information in tweets. V-oc = Valence Ordinal Classification, E-c = Emotion Classification, and E-i = Emotion Intensity Regression.
Framework for Analyzing Emotion
In this section, we present our framework for analyzing emotion. We first establish a basic understanding of emotion polarity by determining the sentiment valence of each root tweet and comment. We then use multi-label emotion detection to predict the emotion categories associated with each post. Based on this data, we explore the interactive nature of emotions, by identifying common patterns in emotion transition pairs between temporally-adjacent posts. Finally we investigate the emotional trajectory within threads to understand how emotional intensity and type shift over time, by aggregating the predicted labels for posts at each time stamp in a given thread. As part of this, we contrast rumour with non-rumour threads, to gain a holistic understanding of emotional expression in rumours and non-rumours on Twitter.
Affective Computing: Automatic Emotion Detection
Manually annotating emotions is both costly and time-consuming, so we use an LLM-based emotion detection model, EmoLLM (Liu et al. 2024a), which is specifically designed for sentiment analysis and emotion detection. The model was instruction-tuned on SemEval 2018 Task1 using a comprehensive emotion labeling scheme grounded in established theoretical frameworks. We prompt the model to perform Valence Ordinal Classification (V-oc), Emotion Classification (E-c), and Emotion Intensity regression (E-i). Detailed prompts are shown in Table 1.
Categorical Emotion Labeling Scheme
Numerous emotion label sets have been proposed (Ekman 1992; Plutchik 1980; Russell 1980). According to Ekman (1992); Plutchik (1980), certain emotions, such as joy, fear, and sadness, are considered more fundamental than others, both physiologically and cognitively. The Valence-Arousal-Dominance (VAD) model (Russell 1980) categorizes emotions within a three-dimensional space of valence (positivity-negativity), arousal (active-passive), and dominance (dominant-submissive). Inspired by Mohammad et al. (2018), we incorporate elements from both basic emotion theories and the VAD model, and further ground EmoLLM emotion predictions to develop the following emotion label schemes: (1) neutral or no emotion; (2) negative emotions: anger (also includes annoyance and rage), disgust (also includes disinterest, dislike, and loathing), fear (also includes apprehension, anxiety, and terror), pessimism (also includes cynicism, and no confidence), sadness (also includes pensiveness and grief); 3) positive emotions: joy (also includes serenity and ecstasy), love (also includes affection), optimism (also includes hopefulness and confidence), anticipation (also includes interest and vigilance), surprise (also includes distraction and amazement) and trust (also includes acceptance, liking, and admiration).
Emotion Polarity: Sentiment Valence
To understand the basic emotion polarity expressed in rumour and non-rumour content, we begin with sentiment valence analysis. Sentiment valence aims to capture the overall emotional tone conveyed by a post, in terms of how positive or negative it is (Liu et al. 2024b). We frame the sentiment valence task as ordinal regression (Mohammad et al. 2018). As shown in Table 1, for a given tweet post, we classify it into one of seven ordinal levels of sentiment intensity, spanning varying degrees of positive and negative valence, that best represents the tweeter’s mental state. The tweet posts within a thread can be divided into two categories: root tweets, which are posted by the publisher, and follow posts, which include all subsequent replies under the root post. We begin by conducting sentiment valence analysis on each post within the thread conversation. For each category, we compute the mean sentiment valence to enable further investigation into the specific emotions associated with different sentiment valences over a thread.
Emotion Distribution
Following sentiment valence analysis, we then examine specific emotions and their distribution in rumour and non-rumour tweet posts. Motivated by the fact that a certain tweet might exhibit more than one emotion, we frame the task as multi-label emotion detection problem. As shown as V-oc in Table 1, given a tweet, we classify it into one of seven ordinal classes, corresponding to various levels of positive and negative sentiment intensity. To reduce noise from automatic emotion detectors, we take the top-three predicted emotions for each tweet. We then aggregate and plot the emotion distribution to provide an overview of dominant emotional trends across the rumour and non-rumour posts. Given that the follow posts make up the majority of the data compared to the root posts, we will focus on using follow posts in our next analysis.
Emotion Transitions
Emotions are contagious and highly interactive (Ferrara and Yang 2015). When publishers write tweets that convey their emotions, readers are likely to respond with emotional reactions of their own (Ferrara and Yang 2015; Zollo et al. 2015). In this part, we model this interactive nature of emotions in the form of emotion transition pairs, which are built from two chronologically-adjacent tweets. In each pair, the first element represents the emotion inferred from the initial content published at a given time, and the second element represents the emotion inferred from the reply content published immediately after. For example, if the first tweet exhibits joy trust and anticipation, and the second tweet shows anger, disgust and surprise, we form the pairs (joy, anger), (joy, disgust), (joy, surprise), (trust, surprise), (trust, surprise), (trust, disgust), (anticipation, anger), (anticipation, surprise) and (anticipation, disgust). We create transitions for all combinations of emotion pairs and explore the likelihood of emotion transition pairs occurring in rumour and non-rumour content. Exploring emotion transitions allows us to understand the emotional flow in social media conversations and uncover typical patterns of rumour and non-rumour content, and any differences between the two.
Emotion Trajectories
We explore the cumulative trajectory of emotion over time to observe how emotions evolve during the conversational thread. We collect all detected emotion labels for each tweet from both rumour and non-rumour content, then track cumulative emotion counts at each chronological step. Finally, we visualize these trends and apply regression models to analyze the growth of emotions over time. This temporal analysis reveals how emotions accumulate or intensify across time, offering insight into the trajectory of emotions in rumour and non-rumour content.
Setting Ru Non p #Ru/Non T F U $p$ (U vs T) $p$ (U vs F) #T/#F/#U
PHEME root --0.25 --0.17 0.00 2602/2602 --0.21 --0.11 --0.39 7.75e-11 4.41e-11 629/629/629
PHEME follow --0.33 --0.26 6.47e-09 --0.35 --0.20 --0.39 0.03 8.38e-15
Twitter15 root --0.26 --0.01 3.51e-05 372/372 --0.21 --0.20 --0.34 0.01 0.01 359/359/359
Twitter15 follow --0.27 --0.06 1.65e-09 --0.24 --0.25 --0.30 0.16 0.21
Twitter16 root --0.18 --0.07 0.00 205/205 --0.11 --0.22 --0.30 1.35e-06 0.18 63/63/63
Twitter16 follow --0.31 --0.12 9.19e-06 --0.30 --0.36 --0.27 0.67 0.90
Table 2: Valence Ordinal Regression results for all datasets. root = root posts, follow = follow posts, Ru = rumour, Non = Non-rumour, T = True rumour, F = False rumour, U = Unverified rumour; $p$ values indicates significance of a one-tailed t-test.
Algorithm 1 PC Algorithm
1:Input: Data $\mathbf{X}$, significance level $\alpha$
2:Output: Completed Partially Directed Acyclic Graph (CPDAG)
3:Initialize a complete undirected graph $G$ with all variables as nodes.
4:Step 1: Skeleton Identification
5:for each pair of variables $(X,Y)$ in $G$ do
6:     Find the subset $S\subseteq\text{Adj}(X,G)\setminus\{Y\}$ such that $X\perp\!\!\!\perp Y\mid S$ with significance $\alpha$.
7:     if such a subset $S$ exists then
8:         Remove the edge $X-Y$ from $G$.
9:     end if
10:end for
11:Step 2: Edge Orientation
12:for each triple of variables $(X,Y,Z)$ in $G$ where $X-Z-Y$ and $X,Y$ are not adjacent do
13:     if $Z\notin S$ for all separating sets $S$ for $X$ and $Y$ then
14:         Orient as $X\to Z\leftarrow Y$ (identify a collider).
15:     end if
16:end for
17:while possible do
18:     for each edge $(X-Y)$ in $G$ do
19:         if there exists a directed path $X\to\dots\to Z$ such that $Z-Y$ then
20:              Orient as $X\to Y$ (acyclicity rule).
21:         else if orienting $X-Y$ as $X\to Y$ creates a new v-structure then
22:              Orient as $X\to Y$ (v-structure rule).
23:         end if
24:     end for
25:end while
26:return the CPDAG representing the equivalence class of causal graphs.
Causal Relationship of Emotions in Rumour & Non-Rumour Threads
To gain a deeper insight into the relationship between rumours and the emotions underlying them, we extend our analysis beyond statistical correlation by conducting a causal analysis. Specifically, we apply the Peter-Clark (PC) algorithm (Spirtes, Glymour, and Scheines 2000), a classical constraint-based causal discovery algorithm on the three merged datasets.
Uncovering causal relations between variables of interest is never an easy problem. Under the fundamental assumption of causal Markov condition that a variable is conditionally independent of all its non-effects given its direct cause, faithfulness ensures that the casual graph exactly encodes the independence and conditional independence relations among variables. These two assumptions allow us to infer causal relationships from observed statistical independencies, forming the cornerstone of constraint-based causal discovery methods.
The PC algorithm identifies causal relationships among the variables of interest, represented as a directed acyclic graph (DAG), by numerating the independence and conditional independence relationships. The algorithm consists of two main steps:
  1. 1.
    Skeleton Identification: Starting with a complete undirected graph where all variables are connected, edges are iteratively removed based on conditional independence and independence relationships among variables, inferred by a conditional independence test. This step returns an undirected graph, which we call a skeleton.
  2. 2.
    Edge Orientation: After constructing the skeleton, edges are oriented by a set of predefined rules (Meek’s Rule (Meek 1997)) to avoid cycles and orient collider structures.
The complete PC algorithm is provided in algorithm 1. It returns a completed partially directed acyclic graph (CPDAG), which represents an equivalence class of causal graphs that are consistent with the observed data’s independence and conditional independence relations. In our implementation, we adopt the Fisher-z test (Fisher 1921) to infer the conditional independence relations.
Results and Discussion
In this section, we apply our framework to the collected data, and present experiment results and discuss our findings.
Figure 1: PHEME Comment Emotion Distribution
Figure 2: Twitter15 Comment Emotion Distribution
Figure 3: Twitter16 Comment Emotion Distribution
Emotion Polarity: Sentiment Valence
We present the sentiment valence ordinal regression results in Table 2. The numbers are balanced by random down-sampling, i.e. rumour and non-rumour, true rumour and false rumour both have equal numbers of posts. As shown in the table, sentiment in rumour root posts and comments is significantly more negative than that in non-rumours across all datasets and settings ($p<0.05$). This means both publishers and commenters engaged in the thread exhibit a more negative mindset towards rumour content. Compared with rumour posts at the root level, comment posts exhibit more negative sentiment for all datasets. Additionally, we break down the rumour data into true, false, and unverified rumours according to their original labels in the dataset. Interestingly, we found that unverified content exhibits more negative sentiment compared to both true and false rumours in the PHEME dataset, as well as in the root posts of Twitter15 and the U vs. T setting in Twitter16. Given that sentiment is more negative in comments and they form the main part of the conversation, we conduct the following experiments using only comments.
Dataset Type Neg Emo Neutral Pos Emo
PHEME Ru 5894 1842 2731
Non 5650 1864 3728
Twitter15 Ru 785 313 636
Non 674 349 776
Twitter16 Ru 438 176 300
Non 388 191 406
Table 3: Statistics of Negative (Neg Emo), Neutral, and Positive Emotions (Pos Emo) across the different datasets for rumour (Ru) and Non-rumour (Non) threads.
Emotion Distribution
In order to further understand emotions expressed in rumour and non-rumour content, we present the emotion distribution results in Figures 1, 2 and 3. Overall, we observe a sharper distribution in emotions like anger, disgust, neutral, optimism, and joy. In the PHEME, Twitter15, and Twitter16 datasets, rumour posts tend to show more negative emotions such as anger, disgust, fear, and sadness, while non-rumour posts display more positive emotions like trust, optimism, joy and love. We present emotions statistics in Table 3.
Figure 4: PHEME rumour emotion transition matrix
Figure 5: Twitter15 rumour emotion transition matrix
Figure 6: Twitter16 rumour emotion transition matrix
Emo Transit PHEME Twitter15 Twitter16
Neg $\rightarrow$ Neg 0.05 0.05 0.07
Neu $\rightarrow$ Neg 0.02 0.06 0.01
Pos $\rightarrow$ Neg 0.07 0.06 0.24
Neg $\rightarrow$ Pos -0.18 0.02 -0.14
Neu $\rightarrow$ Pos -0.13 0.01 -0.08
Pos $\rightarrow$ Pos -0.19 0.02 -0.38
Neg $\rightarrow$ Neu 0.03 -0.20 0.03
Neu $\rightarrow$ Neu 0.08 -0.16 0.07
Pos $\rightarrow$ Neu 0.13 -0.19 -0.05
Table 4: Emotion Transition Delta values across datasets. Emo Transit represents transitions between emotional states, and shows the corresponding delta values for each dataset. Positive values indicate that the pattern occurs more frequently in rumour comments, while negative values mean they are more common in non-rumour comments.
Emo PHEME Twitter15 Twitter16
Ru Non Ru Non Ru Non
Ang 320.50 342.31 41.18 41.75 32.92 25.33
Disg 401.39 420.20 58.67 53.42 44.05 33.77
Fear 154.48 128.42 13.95 7.01 6.56 4.82
Sad 227.86 195.43 45.02 31.97 18.15 18.01
Pess 49.04 37.88 5.26 5.28 3.06 4.32
Neu 408.02 378.33 63.69 75.41 50.41 49.62
Surp 38.79 30.30 14.94 9.52 5.48 5.84
Antic 77.45 87.14 14.42 13.34 8.55 7.98
Trust 26.85 33.94 4.16 5.45 2.89 2.59
Opti 124.37 178.74 24.80 32.19 15.52 20.10
Joy 145.83 195.16 47.65 51.07 24.75 33.98
Love 29.69 55.08 10.97 12.88 4.19 10.16
Table 5: Cumulative emotion regression coefficient across different datasets for rumour and non-rumour comments. Ang = Anger, Disg = Disgutst, Sad = Sadness, Pess = Pessimism, Neu = Neutral, Surp = Surprise, Antic = Anticipation, Opti = Optimism.
Emotion Transitions
We present emotion transition results for each dataset in Figures 4, 5 and 6. The computation was conducted as follows: for each emotion transition pair, we compute the probability based on pair frequency. In order to better reveal the gap between rumours and non-rumours, we define the difference of Emotion Transition (ET) probability as follows:
Emotion Transition (ET): Let’s assume there are $N$ emotions ($N=12$ in our case), let $ET(i,j)$ represent the probability of transitioning from emotion $i$ (i.e. joy) to emotion $j$ (i.e. anger), where $0\leq i<N$ and $0\leq j<N$. This probability is calculated based on the frequency of all pairs that starts with emotion $i$.
$ET(i,j)=\frac{Freq(i,j)}{\Sigma_{k}^{N}Freq(i,k)}$ (1)
Emotion Transition Delta ($\Delta ET$) Define $\Delta ET(i,j)$ as the difference in emotion transition probabilities between rumours and non-rumours for the pair $(i,j)$:
$\Delta ET(i,j)=\frac{ET_{\text{rumour}}(i,j)-ET_{\text{non-rumour}}(i,j)}{ET_{\text{rumour}}(i,j)}$ (2)
Then we visualize it using a heatmap, e.g. in Figure 5, the cell 0.55 in the last row of the third column has is dark red in color, indicating that the emotion transition pair (love, fear) appears more frequently in rumour than non-rumour comments in Twitter15. Overall, we observe larger emotion transition probability mass in positive–positive and negative–negative emotion transitions. This indicates that emotions are contagious, aligning with psychological findings (Goldenberg and Gross 2019; Herrando and Constantinides 2021). Contrasting rumour and non-rumour comments, we observe common patterns, namely that fear–fear and love–sadness are more common in rumour comments, and love–joy and love–optimism appear more frequently in non-rumour comments. We also see differences among datasets: Twitter15 has more anger response to almost all emotions more in non-rumour posts; Twitter16 has a lot of anger and disgust in response to positive emotions in rumours. We aggregate emotions into Negative, Neutral and Positive emotions in Table 4. We observe positive values in emotion pairs where the transition ends with a negative emotion, indicating that discussions in rumours often trigger negative responses. On the contrary, negative delta values are observed in PHEME, Twitter16, suggesting non-rumours tend to prompt more positive responses.
Figure 7: Cumulative Emotion Trajectory of PHEME.
Figure 8: Cumulative Emotion Trajectory of Twitter15.
Figure 9: Cumulative Emotion Trajectory of Twitter16.
Emotion Trajectory
Figures 7, 8 and 9 illustrate the cumulative emotion across each dataset over time. At each chronological step, the counts represent the total number of observed emotions. Generally, we see a strong linear trend across datasets for all emotions. To better capture the rate of growth for each emotion, we apply linear regression and present the slopes in Table 5. From the table, it is apparent that negative emotions tend to grow faster in rumour posts than in non-rumour posts across all datasets, while positive emotions grow faster in non-rumour posts.
Causal Analysis
The causal relationships revealed in Figure 10 demonstrate several key patterns. First, we find out that the fact that a given thread is a rumour is not directly connected with other emotions. Specifically, the rumour has to rely on the emotion of surprise as a bridge to interact with other emotions, namely, $rumour\not\perp\!\!\!\perp Fear|Surprise$ and $rumour\not\perp\!\!\!\perp Anticipation|Surprise$. The change in the distribution of the rumour does not influence other emotions except surprise. This aligns with cognitive basis of rumour transmission, where surprising or counterintuitive information tends to capture attention and facilitate rumour spreading (Knapp 1944; Āllport and Postman 1947). Second, pessimism is primarily influenced by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust). Notably, there’s undirected edge between anger and disgust, this relationship aligns with our previous findings that both rumour and non-rumour posts exhibit intense expressions of these emotions.
Figure 10: Causal graph of is_rumour and emotions. Arrows represents the causal relationships. Orange emotions represent positive emotions, green emotions are negative emotions. Surprise serves as a bridge between is_rumour and other emotions in this context and is depicted in light blue.
There are also a few counterintuitive findings, including sadness leading to joy, joy causing pessimism, and the causal relationship between disgust and love. We conducted a qualitative analysis of the 50 samples and found there are several possible reasons for this: (1) there is a complex interplay of emotions in social media interactions, where emotional responses are shaped by context and individual perspectives. For example, we had one response where joy was detected, “this tweet gives me hope that she may write an eighth” to the post “once again, jk rowling is not working on an eighth harry potter book.” where sadness is detected. Sadness can sometimes lead to joy when people use humor or shared memories to find solace in sadness, which serves as a coping mechanism for processing uncomfortable or shocking topics. Similarly, expressions of joy can paradoxically evoke pessimism in certain contexts, as the same post can be interpreted in vastly different ways depending on the readers’ emotional state, cultural background, or personal experiences. The dual causal relationship between disgust and love further emphasizes the complexity of emotions expressed in text. A post that initially provokes disgust might also elicit admiration or affection when audiences recognize an underlying message of authenticity, vulnerability, or humor. (2) Moreover, the analysis reveals challenges in accurately interpreting emotions through automatic labeling methods. Sarcasm and humor are frequently misclassified, with sarcasm often mistaken for joy due to its seemingly optimistic wording. The lack of contextual information leads to noise and inaccuracies in emotional categorizations. (3) The complexity of social media participants also contributes to this. Some users engage with posts for self-serving purposes, such as promoting their brand or gaining visibility, rather than genuinely responding to the content. These interactions together add a layer of noise to the results, making it even more challenging.
Related Work
The definition of rumour is generally complicated and varies from one publication to another. Some early work treated rumour as information that is false (Cai, Wu, and Lv 2014). Most recent definitions of rumours are “unverified and instrumentally relevant information statements in circulation” (DiFonzo and Bordia 2007) and “unverified information at the time of the posting”. This definition also aligns with the concept in recent work (Zubiaga et al. 2018, 2015; Tian, Zhang, and Lau 2022) and the Oxford English Dictionary, which defines the rumour as “an unverified or unconfirmed statement or report circulating in a community”. [^2] https://www.oed.com/dictionary/rumour_n?tab=meaning_and_use
Existing research highlights the significant role of emotions in understanding general misinformation, mostly fake news. Research has found relationships exist between negative sentiment and fake news, and between positive sentiment and genuine news (Zaeem et al. 2020). Fake news also expresses a higher level of overall emotion, negative emotion, and anger than real news (Zhou, Tao, and Zhang 2022). Negative emotions like sadness and anger can serve as indicators of misinformation (Prabhala and Bose 2019). The role of emotions in rumours has been recognized since the Second World War, reflecting the interactive and community-driven nature of rumour spreading. Knapp’s taxonomy (Knapp 1944) of rumours categorizes them into three types, each deeply embedded with emotions: (1) ‘pipedream’ rumours, which evoke wishful thinking; (2) ‘bogy’ rumours, which heighten anxiety or fear; and (3) ‘wedge-driving’ rumours, which incite hatred. This taxonomy underscores how rumours are inherently embedded with emotional undercurrents.”
Recent research on emotion in rumours largely focuses on their role in spreading behaviour, some studies have used questionnaires to gather participants’ reactions to specific rumours (Zhang et al. 2022; Rijo and Waldzus 2023; Ali et al. 2022), while others have employed cascade size and lifespan as indicators (Pröllochs, Bär, and Feuerriegel 2021b, a). Key findings of such work include: rumours conveying anticipation, anger, trust, or offensiveness tend to generate more shares, have longer lifespans, and exhibit higher virality (Pröllochs, Bär, and Feuerriegel 2021b). Additionally, false rumours containing a high proportion of terms reflecting positive sentiment, trust, anticipation, anger, or condemnation are more likely to go viral (Solovev and Pröllochs 2022; Pröllochs, Bär, and Feuerriegel 2021a). However, existing research has notable gaps: it often focuses on isolated aspects of emotions in rumours, primarily identifies correlations rather than causality, and tends to examine rumour data in isolation. To address these gaps, we aim to propose a comprehensive emotional analytical framework that integrates multiple emotion-related tasks, contrasting rumour and non-rumour content with the aim to enhance our understanding of emotions and, ultimately, to improve rumour detection in online social media.
Conclusion
In this work, we presented a analytical emotion framework for online rumours. We make the use of EmoLLM for automatic emotion detection and ground its predictions to the categorical emotion labeling scheme. The framework analyzes the emotion from various aspects, from simply emotion polarity (sentiment valence), emotion distribution, emotion transition and trajectory, we also analyze the causal relationship between rumour label and emotions. The key findings include: compared with non-rumour contents, rumour are significantly more negative in sentiments, containing more negative emotions like anger, fear and pessimism; emotions are contagious in online social context, rumour contents usually trigger negative responses and non-rumours tend to receive positive ones; Cumulative emotion regression coefficient showed that negative emotions grow significantly faster in rumours comments as positive emotions in non-rumour ones; the rumour tweets are not directly connected with other emotions and rely on the emotion surprise as a bridge. Pessimism is primarily influenced by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust), anger and disgust exhibit bidirectional causation. By presenting the framework, we hope to facilitate research in more comprehensive and fine-grained study in emotion in online rumour contents and better detection techniques.
Limitations and future work
This work also faces several challenges and limitations: 1) We rely on EmoLLM as our automatic emotion detection tool for all emotion-related tasks. While it is generally efficient and effective, it exhibits inaccuracies in analyzing complex online discussions, such as those involving sarcasm; 2) Although we have access to the chronological order of tweets within conversations, explicit conversation structures are not available for all data; and 3) The datasets used in this study are limited to English textual rumour data. Future work should explore multilingual and multimodal content in rumour conversations to provide a more comprehensive analysis.
Ethical Impacts
Analyzing emotions in rumour detection presents ethical challenges, such as privacy invasion, interpretative biases, risks of emotional manipulation, amplification of harmful content, and cultural insensitivity. To address these concerns, we advocate for responsible and transparent use, prioritizing individual privacy and freedom of expression, with clear communication and opt-out options for users. This research was conducted independently using publicly available datasets, and the framework was developed to enhance academic understanding and combat misinformation online for the public good.
License of Artifacts
We list the licenses of different artifacts used in this paper: PHEME [^3] https://figshare.com/articles/dataset/PHEME_dataset_for_Rumour
_Detection_and_Veracity_Classification/6392078
is under CC-BY license, Twitter15 and Twitter 16444https://github.com/majingCUHK/rumour_RvNN are under MIT License, EmoLLM [^5] https://github.com/lzw108/EmoLLMs is under MIT License and Huggingface Transformers [^6] https://github.com/huggingface/transformers is under Apache License 2.0). Our source code and annotated data will be under MIT license.